# Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap

### v1

@@ 4.1 Evaluating Different Geometric Regularizations
+ (DST) $26.8$$\pm 3.6$ $33.7$$\pm 4.2$ $26.8$$\pm 3.6$ $33.7$$\pm 4.2$
+ (GMG) 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$ 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$
Thus, next we benchmark the GMG in its CR form against its distortion counterpart across three more datasets again over four different baselines. We report full results in Table 2. Again we observe that the GMG outperforms or performs equally well to its distortion equivalent, confirming the benefits of accounting for the optimal possible mapping in the regularization. Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding an HFS regularization and obtained the best results without it. We emphasize that using the GMG as CR significantly improves results for all datasets versus not using any isometric regularization. This establishes the GMG as an effective regularization method beneficial for disentangled representation learning.
Regularizing the Encoder.
Lastly, we also analyze a CR on $e_{\phi}$, as well as regularizing both $d_{\theta}$ and $e_{\phi}$ together. We report full results over two datasets in Table 4. Again, the GMG on $d_{\theta}$ achieves best DCI-D over all baselines. This result is expected in the light of Theorem 3.7. Interestingly, regularizing solely $d_{\theta}$ outperforms regularizing both $e_{\phi}$ and $d_{\theta}$. We hypothesize this is due to the regularization of the decoder also offering a stronger signal as its gradients impact both the decoder and the encoder, as in this case, the reference $r$ is the distribution of encoded images.
4.2 Towards Decoder-free Disentanglement
Recently, works such as [11, 70, 21, 47, 1] have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation
Table 3: Disentanglement (DCI-D) without a decoder trained with various regularizations on Shapes3D [9].
Decoder-free $\beta$-VAE $\beta$-TCVAE Base $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ Isometric (IR) + (DST) $38.2$$\pm 0.8$ $42.7$$\pm 1.6$ + (GMG) $13.9$$\pm 0.4$ $20.5$$\pm 0.5$ Scaled Isometric (SIR) + (DST) 45.6 ±$\pm 1.2$ 53.5 ±$\pm 1.0$ + (GMG) $15.2$$\pm 0.3$ $25.2$$\pm 0.6$ Conformal (CR) + (DST) $37.0$$\pm 0.4$ $46.1$$\pm 1.5$ + (GMG) $37.0$$\pm 0.9$ $38.8$$\pm 1.1$
learning [14, 72, 3, 24]. However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.
@@ 5 Conclusion
In this work, we introduce an optimal transport (OT) perspective on unsupervised disentangled representation learning to incorporate general latent geometrical constraints. We derive the Gromov-Monge gap (GMG), a provably weakly convex OT regularizer that measures the preservation of geometrical properties by a transport map $T$. By formulating disentangled representation learning as a transport problem, we integrate the GMG into standard training objectives, allowing for incorporating and studying various geometric constraints on the disentanglement of learned representation spaces. Including these geometry preserving regularization offers significant performance benefits across four standard disentanglement benchmarks when applied to existing disentanglement methods. Moreover, we show promising results on decoder-free unsupervised disentanglement. We demonstrate that optimizing for geometric constraints through the OT lens can provide sufficient training signal and regularization on the model encoder to achieve measurable disentanglement without explicit reconstruction objectives. This opens a possible door towards more scalable unsupervised disentanglement and bridges to weakly- & self-supervised encoder-only representation learning efforts.
@@ Acknowledgements
Appendix A Proofs
A.1 Positivity of the Entropic GMG
Recall that
$\begin{split}\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}(T)&\vcentcolon=\tfrac{1}{n}\mathcal{D}_{r_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,\\ &=\mathcal{D}_{r_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,\end{split}$
As a result, applying minimization on both sides yields that $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({r}_{n},T\sharp{r}_{n})<\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}({r}_{n},T\sharp{r}_{n})$, and therefore:
$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}(T)>\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}(T)=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\geq 0.$
A.2 Reminders on Monge and Kantorovich OT
@@
Instead of intra-domain cost functions, we consider here an inter-domain continuous cost function $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. This assumes that we have a meaningful way to compare elements $\mathbf{x},\mathbf{y}$ from the source and target domains. The Monge [49] problem (MP) between $p\in\mathcal{P}(\mathcal{X})$ and $p\in\mathcal{P}(\mathcal{Y})$ consists of finding a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $p$, while minimizing the average displacement cost quantified by $c$
@@
$\mathrm{W}_{c}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}c(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\,.$ (KP)
@@
A.3 Conditionally Positive Kernels
@@
A kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is conditionally positive if it is symmetric and for any $\mathbf{x}_{1},...,\mathbf{x}_{n}\in\mathbb{R}^{d}$ and $\mathbf{a}\in\mathbb{R}^{n}$ s.t. $\mathbf{a}^{\top}\mathbf{1}_{n}=0$, one has
@@
Conditionally positive kernels include all positive kernels, such as the inner-product $k(\mathbf{x},\mathbf{y})=\langle\mathbf{x},\mathbf{y}\rangle$, the cosine similarity $k(\mathbf{x},\mathbf{y})=\textrm{cos-sim}(\mathbf{x},\mathbf{y})=\langle\tfrac{\mathbf{x}}{\|\mathbf{x}\|_{2}},\tfrac{\mathbf{y}}{\|\mathbf{y}\|_{2}}\rangle$, but also the negative squared Euclidean distance $k(\mathbf{x},\mathbf{y})=-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. Therefore, each of the costs of interest is either a conditionally positive kernel - for the inner product and the cosine distance - or its opposite is - the squared Euclidean distance.
Appendix B Proofs of § 3.2
@@
Let $T,r,s$ as described and suppose that $\mathcal{GM}_{r}^{c}(T)=0$. Then, $\pi^{r}:=(\mathrm{Id},T)\sharp r$ is an optimal Gromov-Wasserstein coupling, solution of Problem (GWP) between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$. Therefore, from [65, Theorem. 3], $\pi^{r}$ is an optimal Kantorvich coupling, solution of Problem (KP) between $r$ and $T\sharp r$ for the linearized cost:
@@
Afterwards, by virtue of [61, Theorem 1.38], $\mathrm{Spt}\left(\pi^{r}\right)$ is a $\tilde{c}$-cyclically monotone (CM) set (see [61, Definition. 1.36]). From the definition of cyclical monotonicity, this property translates to subsets. Then, by defining $\pi^{s}=(\mathrm{Id},T)\sharp s$, as $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$, one has $\mathrm{Spt}(\pi^{s})=\mathrm{Spt}((\mathrm{Id},T)\sharp s)\subset\mathrm{Spt}((\mathrm{Id},T)\sharp r)=\mathrm{Spt}(\pi^{r})$, so $\mathrm{Spt}(\pi^{s})$ is $\tilde{c}$-CM. Finally, since $\mathcal{X}$ and $\mathcal{Y}$ are compact, and $\tilde{c}$ is uniformly continuous, the $\tilde{c}$-cyclical monotonicity of its support implies that the coupling $\pi^{p}$ is a Kantorovich optimal coupling between its marginals for cost $\tilde{c}$, thanks to [61, Theorem 1.49]. By re-applying [65, Theorem. 3], we get that $\pi^{s}$ solves the Gromov-Wasserstein problem between its marginals for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$. In other words, $\pi^{s}=(\mathrm{Id},T)\sharp s$ is Gromov-Wasserstein optimal coupling between $s$ and $T\sharp s$ so $T$ is a Gromov-Monge map between $s$ and $T\sharp s$ and $\mathcal{GM}_{s}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=0$. ∎
Appendix C Proofs of § 3.3
See 3.4
@@
We start by showing a more general results, stating that when $c_{\mathcal{X}},c_{\mathcal{Y}}$ are conditionally positive kernels (see A.1), the discrete GW couplings between uniform, empirical distributions supported on the same number of points, ae permutation matrices.
Proposition C.1 (Equivalence between Gromov-Monge and Gromov-Wasserstein problems in the discrete case.).
@@
$\begin{split}\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({p}_{n},{p}_{n})&=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\sigma\in\mathcal{S}_{n}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{\sigma(i)},\mathbf{y}_{\sigma(j)}))^{2}\end{split}$ (9)
@@
As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels, from [65, Theorem. 3], $\mathbf{P}^{\star}$ also solves:

### v2

@@ 4.2 Leveraging the Gromov-Monge Gap for Disentanglement
+ Dec-DST $26.8$$\pm 3.6$ $33.7$$\pm 4.2$ $26.8$$\pm 3.6$ $33.7$$\pm 4.2$
+ Dec-GMG 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$ 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$
GMG Consistently Enhances Disentanglement.  To further validate our findings, we benchmark the GMG for decoder regularization with angle preservation ($\mathbf{Cos}$) against its distortion counterpart across three more datasets. We report full results in Table 2. Again we observe that the GMG outperforms or performs equally well to its distortion equivalent, confirming the benefits of accounting for the optimal possible mapping in the regularization. Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding an HFS regularization and obtained the best results without it. We emphasize that using the GMG with $\mathbf{Cos}$ regularization significantly improves results for all datasets versus not using any isometric regularization. This establishes the GMG as an effective tool for enhanced disentanglement.
Figure 3: Stability analysis for the DST and the GMG applied to the decoder $d_{\theta}$ with cost $\mathbf{Cos}$. For each regularizer, we assess stability by measuring the alignment of its gradients across 5 batches. We compute the cosine similarity between all pairs of gradients. Values closer to 1 indicates better alignment, hence better stability.
Stability of the GMG. Furthermore, we investigate the stability of the GMG compared to the DST. Computing the GMG using Algorithm 1 involves solving an optimization problem through a GW solver. This experiment aims to demonstrate that our proposed method for solving the GW problem leads to a stable loss in the GMG . We assess the stability by measuring the alignment of GMG gradients $\nabla_{\theta}\mathrm{GMG}_{r_{n}}(T_{\theta})$ across 5 randomly sampled batches from each dataset $r_{n}\sim\mathcal{D}$, for a fixed neural map $T_{\theta}$. We repeat this procedure for each of the four datasets $\mathcal{D}$ considered and apply the same methodology to the DST. Figure 4 presents results for our best-performing configuration, which regularizes the decoder (i.e., $T_{\theta}=d_{\theta}$) and employs the $\mathbf{Cos}$ cost. Results for encoder regularization (i.e., $T_{\theta}=e_{\phi}$) are reported in Appendix D.1. We observe that the GMG’s gradients exhibit significantly higher alignment compared to those of the DST, demonstrating greater stability of our proposed regularizer. This suggests that the GMG ’s formulation, which accounts for the the minimal distortion, effectively mitigates the inherent variability inherent of the distortion, leading to more stable gradient computations.
@@ 5 Conclusion
In this work, we introduce an OT perspective on unsupervised disentangled representation learning to incorporate general latent geometrical constraints. We derive the GMG, a provably weakly convex regularizer that measures whether a map $T$ transports a fixed reference distribution with minimal distortion of some predefined geometric features. By formulating disentangled representation learning as a transport problem, we integrate the GMG into standard training objectives, allowing for incorporating and studying various geometric constraints on the learned representation spaces. We show significant performance benefits of our approach on four standard disentanglement benchmarks.
6 Reproduciblity
In this work, we introduce the GMG, computed as detailed in Algorithm 1. To facilitate reproducibility, we provide the implementation code for computing the GMG on source and target batches in Appendix E. Comprehensive proofs, including all underlying assumptions, are presented in Appendix B. For our experiments on disentanglement benchmarks, we adhere to standard practices, employing streamlined preprocessing across all datasets. Detailed descriptions of these procedures are available in Appendix C. All experiments described in this paper can be conducted using a single RTX 2080TI GPU, ensuring accessibility and replicability of our results.
@@ Acknowledgements
The Appendix is organized as follows:
  • Section A provides additional background information to supplement the main text.
  • Section B presents all theoretical proofs, including detailed assumptions.
  • Section C outlines comprehensive experimental details, ensuring reproducibility.
  • Section D offers supplementary empirical results that further support our findings.
  • Section E includes the implementation code for computing the GMG.
Appendix A Additional Background
A.1 Reminders on Monge and Kantorovich OT
@@
Instead of intra-domain cost functions, we consider here an inter-domain continuous cost function $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. This assumes that we have a meaningful way to compare elements $\mathbf{x},\mathbf{y}$ from the source and target domains. The Monge (1781) problem (MP) between $p\in\mathcal{P}(\mathcal{X})$ and $p\in\mathcal{P}(\mathcal{Y})$ consists of finding a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $p$, while minimizing the average displacement cost quantified by $c$
@@
$\mathrm{W}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}c(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\,.$ (KP)
@@
A.2 Conditionally Positive Kernels
@@
A kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is CPD, i.e., conditionally positive, if it is symmetric and for any $\mathbf{x}_{1},...,\mathbf{x}_{n}\in\mathbb{R}^{d}$ and $\mathbf{a}\in\mathbb{R}^{n}$ s.t. $\mathbf{a}^{\top}\mathbf{1}_{n}=0$, one has
@@
CPD include all positive kernels, such as the inner-product $k(\mathbf{x},\mathbf{x}^{\prime})=\langle\mathbf{x},\mathbf{x}^{\prime}\rangle$, or the cosine similarity $k(\mathbf{x},\mathbf{x}^{\prime})=\textrm{cos-sim}(\mathbf{x},\mathbf{x}^{\prime})=\langle\tfrac{\mathbf{x}}{\|\mathbf{x}\|_{2}},\tfrac{\mathbf{x}^{\prime}}{\|\mathbf{x}^{\prime}\|_{2}}\rangle$, but also the (scaled) negative squared Euclidean distance $k(\mathbf{x},\mathbf{x}^{\prime})=-\alpha\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}$, $\alpha>0$. Therefore, each of the costs of interest is either a conditionally positive kernel - for the inner product and the cosine distance - or its opposite is - for the squared Euclidean distance. Additionally, CPD kernels also include more exotic cost functions, such as:
  • The RBF kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}/\gamma),\gamma>0$.
  • The power kernel $k(\mathbf{x},\mathbf{x}^{\prime})=-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{p},0<p<2$.
  • The thin plate spline kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}\log(\|\mathbf{x}-\mathbf{x}^{\prime})$
  • The inverse multi-quadratic kernel $k(\mathbf{x},\mathbf{x}^{\prime})=1/\sqrt{\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}+c^{2}},c\in\mathbb{R}$.
As a result, the family of CPD kernels includes a large variety of cost functions, which can be used to define various GMGs.
Appendix B Proofs
B.1 The GMG characterizes Gromov-Monge Optimality
We show here that if $\mathrm{GMG}_{r}(T)=0$, then $T$ is a Gromov-Monge map between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. As the set of deterministic couplings $\{\pi_{F}\vcentcolon=(\mathrm{I}_{d},F)\sharp r|F:\mathcal{X}\to\mathcal{Y}\,,F\sharp r=T\sharp r\}\subset\Pi(r,T\sharp r\}$, we immediately get that
$\inf_{F\sharp r=T\sharp r}\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(F(\mathbf{x}),F(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\geq\mathrm{GW}(r,T\sharp r)$ (7)
On the other hand, if $\mathrm{GMG}_{r}(T)=0$, one has
$\mathrm{GW}(r,T\sharp r)=\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})$ (8)
Therefore, combining Eq. (7) and Eq. (8), we get
$\begin{split}&\inf_{F\sharp r=T\sharp r}\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(F(\mathbf{x}),F(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\\ =&\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\end{split}$ (9)
Finally, as $T$ naturally satisfies the marginal constraint, we conclude that $T$ is a Gromov-Monge map between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$.
B.2 On rescaling the costs matrices in the entropic GW solver
We remind, from Eq. (EGWP), that
$\mathrm{GW}_{\varepsilon}({p}_{n},{q}_{n})=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(\mathbf{C}_{\mathcal{X}_{i,i^{\prime}}}-\mathbf{C}_{\mathcal{Y}_{j,j^{\prime}}})^{2}\,\mathbf{P}_{i,j}\mathbf{P}_{i^{\prime},j^{\prime}}-\varepsilon H(\mathbf{P})\,.$ (10)
By developping each terms, and using the fact that $\mathbf{P}\in U_{n}$, we get
$\begin{split}\mathrm{GW}_{\varepsilon}({p}_{n},{q}_{n})&=\min_{\mathbf{P}\in U_{n}}\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{X}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle+\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{Y}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\\ &=\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{X}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle+\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{Y}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle+\min_{\mathbf{P}\in U_{n}}-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\end{split}$ (11)
where $\mathbf{C}_{\mathcal{X}}^{\odot 2}=\mathbf{C}_{\mathcal{X}}\odot\mathbf{C}_{\mathcal{X}}$, with $\odot$ the Hadamard (i.e., elementwise) product, and similarly for $\mathbf{C}_{\mathcal{Y}}^{\odot 2}$. As we can see that the two terms on the left do not depend on $\mathbf{P}$, they do not impact the minimization, an OT coupling $\mathbf{P}^{\star}$ solving the problem satisfies:
$\begin{split}\mathbf{P}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,.\end{split}$ (12)
As a result, if we now replace $\mathbf{C}_{\mathcal{X}}$ and $\mathbf{C}_{\mathcal{Y}}$ by $\mathbf{C}_{\mathcal{X}}/\texttt{stat}(\mathbf{C}_{\mathcal{X}})$ and $\mathbf{C}_{\mathcal{Y}}/\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$, respectively, the new OT coupling $\tilde{\mathbf{P}}^{\star}$ solving the problem satisfies
$\begin{split}&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\frac{\mathbf{C}_{\mathcal{X}}}{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\mathbf{P}\frac{\mathbf{C}_{\mathcal{Y}}}{\texttt{stat}(\mathbf{C}_{\mathcal{Y}})},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-\frac{2}{{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})}\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})\cdot\varepsilon H(\mathbf{P})\,,\end{split}$ (13)
where in the last line, we use the fact that $\texttt{stat}(\mathbf{C}_{\mathcal{X}}),\texttt{stat}(\mathbf{C}_{\mathcal{Y}})>0$. This yields the desired equivalence on scaling the cost matrix and adapting the entropic regularization strength.
B.3 Positivity of the Entropic GMG estimator
Recall that
$\begin{split}\mathrm{GMG}_{{r}_{n},\varepsilon}(T)&=\mathrm{DST}_{r_{n}}(T)-\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,\\ &=\mathrm{DST}_{r_{n}}(T)-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,\end{split}$
As a result, applying minimization on both sides yields that $\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})<\mathrm{GW}_{0}({r}_{n},T\sharp{r}_{n})=\mathrm{GW}({r}_{n},T\sharp{r}_{n})$, and therefore:
$\mathrm{GMG}_{r_{n},\varepsilon}(T)>\mathrm{GMG}_{r_{n},0}(T)=\mathrm{GMG}_{r_{n}}(T)\geq 0.$
B.4 Proofs of Prop. 3.3
@@ B.3 Positivity of the Entropic GMG estimator
Let $T,r,s$ as described and suppose that $\mathcal{GM}_{r}^{c}(T)=0$. Then, $\pi^{r}:=(\mathrm{Id},T)\sharp r$ is an optimal Gromov-Wasserstein coupling, solution of Problem (GWP) between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$. Therefore, from (Séjourné et al., 2023, Theorem. 3), $\pi^{r}$ is an optimal Kantorvich coupling, solution of Problem (KP) between $r$ and $T\sharp r$ for the linearized cost:
@@
Afterwards, by virtue of (Santambrogio, 2015, Theorem 1.38), $\mathrm{Spt}\left(\pi^{r}\right)$ is a $\tilde{c}$-cyclically monotone (CM) set (see (Santambrogio, 2015, Definition. 1.36)). From the definition of cyclical monotonicity, this property translates to subsets. Then, by defining $\pi^{s}=(\mathrm{Id},T)\sharp s$, as $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$, one has $\mathrm{Spt}(\pi^{s})=\mathrm{Spt}((\mathrm{Id},T)\sharp s)\subset\mathrm{Spt}((\mathrm{Id},T)\sharp r)=\mathrm{Spt}(\pi^{r})$, so $\mathrm{Spt}(\pi^{s})$ is $\tilde{c}$-CM. Finally, since $\mathcal{X}$ and $\mathcal{Y}$ are compact, and $\tilde{c}$ is uniformly continuous, the $\tilde{c}$-cyclical monotonicity of its support implies that the coupling $\pi^{p}$ is a Kantorovich optimal coupling between its marginals for cost $\tilde{c}$, thanks to (Santambrogio, 2015, Theorem 1.49). By re-applying (Séjourné et al., 2023, Theorem. 3), we get that $\pi^{s}$ solves the Gromov-Wasserstein problem between its marginals for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$. In other words, $\pi^{s}=(\mathrm{Id},T)\sharp s$ is Gromov-Wasserstein optimal coupling between $s$ and $T\sharp s$ so $T$ is a Gromov-Monge map between $s$ and $T\sharp s$ and $\mathrm{GMG}_{s}(T)=0$. ∎
B.5 Proofs of Thm. 3.5
See 3.5
Before proving Thm. 3.5, we first demonstrate some technical results that will be useful later.
Reformulation of the empirical GMG using permutations.
We start by showing that $\mathrm{GMG}_{{r}_{n}}(T)$ is always the sub-optimality gap of $T$ in Prob. (GMP) between $r_{n}$ and $T\sharp r_{n}$. This occurs because Prob. (GMP) and Prob. (GWP) coincide when applied between empirical measures on the same number of points. In other words, we can reformulate Prob. (GWP) between $r_{n}$ and $T\sharp r_{n}$ using permutation matrices, instead of (plain) couplings.
Proposition B.1.
The empirical GMG reads
$\begin{split}\mathrm{GMG}_{{r}_{n}}(T)=\mathrm{DST}_{{r}_{n}}(T)-\min_{\sigma\in\mathcal{S}_{n}}\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\left(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{\sigma(i)}),T(\mathbf{x}_{\sigma(j)}))\right)^{2}\end{split}$ (15)
@@
We first show a more general results, stating that when $c_{\mathcal{X}},c_{\mathcal{Y}}$ are conditionally positive kernels (see A.1), the discrete GW couplings between uniform, empirical distributions supported on the same number of points, ae permutation matrices.
Proposition B.2 (Equivalence between Gromov-Monge and Gromov-Wasserstein problems in the discrete case.).
@@
$\begin{split}\mathrm{GW}({p}_{n},{p}_{n})&=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\sigma\in\mathcal{S}_{n}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{\sigma(i)},\mathbf{y}_{\sigma(j)}))^{2}\end{split}$ (16)
@@
As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels, from (Séjourné et al., 2023, Theorem. 3), $\mathbf{P}^{\star}$ also solves: